Use the elimination method to solve the following system of equations. what is z,y & z
x-2y+z=3
2x+y-2z=-7
3x-y+z=0

Can't seem to get this one right, Thank you!

Answer by TimothyLamb(4379)

x-2y+z=3
2x+y-2z=-7
3x-y+z=0
---
put the system of linear equations into standard form
---
x-2y+z=3
2x+y-2z=-7
3x-y+z=0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -1
y = -1
z = 2
---
now by elimination:
---
2x + y - 2z = -7
3x - y + z = 0
---
add the above two equations to eliminate y = yields equation A (in x and z)
---
x - 2y + z = 3
2(2x + y - 2z = -7)
---
add the above two equations to eliminate y = yields equation B (in x and z)
---
do a linear combination of equations A and B to eliminate x and solve for z = 2
---
a linear combination of equations A and B is generally = N*A + M*B (where N and M are integers)
---
with z = 2 substitute into equation A to get x = -1
---
with z = 2 and x = -1 substitute into any of the original equations and solve for y = -1
---
solution by elimination:
x = -1
y = -1
z = 2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---